Solve the equation ln(x+5)-ln(3)=ln(x-3).
Answer by venugopalramana(3286)

Solve the equation ln(x+5)-ln(3)=ln(x-3).
ln[(x+5)/3]=ln(x-3)
(x+5)/3=(x-3)
x+5=3(x-3)=3x-9
3x-x=5+9=14
2x=14
x=7